# Static Electric Dipole Polarizability and Hyperpolarizability Tensors from Mean-Field Cavity Quantum Electrodynamics Approaches

### v1

@@ Abstract
First-order electric dipole response functions are implemented for cavity quantum electrodynamics (QED) generalizations of Hartree-Fock (HF) and Kohn-Sham density functional theory (DFT) in order to assess the degree to which static molecular response properties are impacted by interactions between electronic degrees of freedom and an optical cavity mode. Isotropically averaged static electric dipole polarizability tensors from QED-HF and QED-DFT are found to be somewhat insensitive to the presence of the cavity under realistic single-molecule coupling strengths. In contrast, the first hyperpolarizability tensor computed using QED-HF can be significantly modified by the cavity, depending on the relative orientation of the molecule and cavity mode polarization axis. For example, compared to the isolated molecule case, the isotropically averaged hyperpolarizability for -nitroaniline decreases by more than 20% when the molecule is coupled to a single-mode optical cavity with a coupling strength of a.u. and when the cavity mode is polarized along the principal molecular axis. On the other hand, with this coupling strength and polarization, the isotropically averaged static dipole polarizability from QED-HF or QED-DFT decreases by only by only 2–5 %, depending on the choice of DFT functional.
@@ IV Results and Discussion
Lastly, the orientation dependence of the polarizability and hyperpolarizability deserve additional discussion. As can be seen in Fig. 1, the isotropically averaged static electric dipole polarizability decreases with increased coupling strength for all cavity mode polarizations we have considered. The story is quite different for the static first hyperpolarizability, where we find that is slightly enhanced when the molecule interacts with and - or -polarized cavity mode, whereas it decreases significantly when interacting with a cavity mode polarized in the -direction. It is well known that cavity induced changes to electronic structure are highly dependent upon the relative orientations of the molecule and cavity mode polarization axis. It has also been argued that meaningful predictions of cavity modified properties should either involve appropriate orientational averaging [84, 83] or account for cavity-induced rotations and geometric distortions.[55] From the geometry relaxation point of view, we note that the energy of -nitroaniline is the lowest when the cavity mode is polarized along the -axis, which happens to be the orientation associated with the second-smallest change to the polarizability and the smallest change to the hyperpolarizability. Hence, our ability to engineer large changes to response properties via cavity interactions depends on our ability to precisely control the relative orientations of the molecule and cavity mode polarization axis.
@@ V Conclusions
A large number of recent computational studies have explored the influence of strong electron–photon coupling on the electronic structure of cavity-embedded molecular systems. Many such studies have focused on the impact that strong coupling can have on ground-state properties, with comparatively fewer works focusing on molecular response properties, particularly higher-than-first-order responses. In this work, we have developed mean-field approaches to model the response of a molecule within an optical cavity to static external electric fields. While the electric dipole polarizability (a first-order response property) is not particularly sensitive to the presence of the cavity, the first hyperpolarizability (a second-order property) can be significantly impacted by the cavity, depending on the cavity mode polarization axis. Moving forward, it will be interesting to examine cavity-induced changes to higher-order static response properties, as well as their dynamic analogues, in which case we expect the cavity mode frequency to play an important part. Lastly, it also will be important to consider how explicit electron–photon correlations that are captured by many-body ab initio QED approaches impact the response properties of cavity-embedded systems, as it has repeatedly been demonstrated that mean-field approaches such as those studied herein tend to exaggerate cavity-induced changes to electronic structure.[69, 54, 55]
@@
The cartesian coordinates for -nitroaniline obtained from geometry optimizations using HF, SVWN, PBE, B3LYP, and B97X and the d-aug-cc-pVTZ basis set, the ground-state energies and response properties of -nitroaniline computed using QED-HF and QED-DFT at different values and a fixed cavity frequency of 0.1 E, the excitation energies and oscillator strenghts of the first bright state of -nitroaniline obtained using QED-HF and QED-DFT, and the ground-state energies and response properties of -nitroaniline computed using QED-HF and QED-DFT at a.u. and different cavity frequencies.
Acknowledgements.

### v2

@@ Abstract
First-order electric dipole response functions are implemented for cavity quantum electrodynamics (QED) generalizations of Hartree-Fock (HF) and Kohn-Sham density functional theory (DFT) in order to assess the degree to which static molecular response properties are impacted by interactions between electronic degrees of freedom and an optical cavity mode. Isotropically averaged static electric dipole polarizability tensors from QED-HF and QED-DFT are found to be somewhat insensitive to the presence of the cavity under realistic single-molecule coupling strengths. In contrast, the first hyperpolarizability tensor computed using QED-HF can be significantly modified by the cavity, depending on the relative orientation of the molecule and cavity mode polarization axis. For example, compared to the isolated molecule case, the isotropically averaged hyperpolarizability for -nitroaniline decreases by more than 20% when the molecule is coupled to a single-mode optical cavity with a coupling strength of a.u. and when the cavity mode is polarized along the principal molecular axis. On the other hand, with this coupling strength and polarization, the isotropically averaged static dipole polarizability from QED-HF or QED-DFT decreases by only 2–5 %, depending on the choice of DFT functional.
@@ III Computational Details
Lastly, we also examine the convergence of the cavity-induced changes in the static polarizability and first hyperpolarizability of -nitroaniline using the cc-pVZ, aug-cc-pVZ, and d-aug-cc-pVZ basis sets with D, T, and Q. These calculations were performed using HF and at the geometry optimized at the HF/d-aug-cc-pVTZ level of theory mentioned earlier. Each calculation was carried out using the default JK-type auxiliary basis sets for cc-pVZ, aug-cc-pVZ, and d-aug-cc-pVZ in Psi4, which are cc-pVZ-JKFIT,[133] aug-cc-pVZ-JKFIT,[133] and def2-universal-JKFIT,[132] respectively.
@@ IV Results and Discussion
The orientation dependence of the polarizability and hyperpolarizability deserve additional discussion. As can be seen in Fig. 1, the isotropically averaged static electric dipole polarizability decreases with increased coupling strength for all cavity mode polarizations we have considered. The story is quite different for the static first hyperpolarizability, where we find that is slightly enhanced when the molecule interacts with and - or -polarized cavity mode, whereas it decreases significantly when interacting with a cavity mode polarized in the -direction. It is well known that cavity induced changes to electronic structure are highly dependent upon the relative orientations of the molecule and cavity mode polarization axis. It has also been argued that meaningful predictions of cavity modified properties should either involve appropriate orientational averaging [84, 83] or account for cavity-induced rotations and geometric distortions.[55] From the geometry relaxation point of view, we note that the energy of -nitroaniline is the lowest when the cavity mode is polarized along the -axis, which happens to be the orientation associated with the second-smallest change to the polarizability and the smallest change to the hyperpolarizability. Hence, our ability to engineer large changes to response properties via cavity interactions depends on our ability to precisely control the relative orientations of the molecule and cavity mode polarization axis.
Lastly, we consider the convergence of the cavity-induced changes in the response properties considered above as a function of the basis set size. Figure 3 illustrates the percent change in the isotropically averaged static electric dipole polarizability as a function of the coupling strength, when the calculations are carried out using the cc-pVZ, aug-cc-pVZ, and d-aug-cc-pVZ basis sets (with D, T, and Q). Cavity-induced changes to the values for -nitroaniline are essentially converged at at any -level, provided that one uses augmented (aug-cc-pVZ) or doubly-augmented (d-aug-cc-pVZ) basis sets. On the other hand, the percent changes in obtained using the cc-pVZ basis sets differ significantly from results obtained from (doubly-) augmented basis sets and also display a clear dependence on the -level. At = 0.05 a.u., the percent change in predicted in these basis sets shows a 0.2% spread, and these values themselves differ by 0.2%–0.4% from the converged d-aug-cc-pVZ results. These observations are similar to those reported for the smaller LiH, H2O, and CO molecules in Ref. 96. Note, also, that calculations carried out in the cc-pVDZ basis set predict that the magnitude of the isotropically averaged polarizability increases with coupling strength when the cavity is polarized in the -direction, which is qualitatively incorrect, relative to results obtained from calculations that use basis sets with diffuse functions.
Figure 4 illustrates the percent change in the isotropically averaged static first hyperpolarizability as a function of the coupling strength, when the calculations are carried out using the same basis sets considered in Fig. 3. As in Fig. 3, the predicted changes in obtained from calculations carried out in non-augmented basis sets are significantly different from those obtained from calculations that use doubly-augmented basis sets; at a.u., cc-pVZ-derived results differ from those obtained using the d-aug-cc-pVQZ basis by roughly 1%–2%, with the results from the z-polarized cavity having the smallest discrepancies. Unlike in Fig. 3, we find that results obtained using aug-cc-pVZ basis sets display a non-trivial -level dependence [see Fig. 4 (B), in particular]. The data included in the Supporting Information indicate that, in order to obtain values that agree to within 1% with those obtained from calculations performed using the largest basis set considered in this work (d-aug-cc-pVQZ), one should use at least an aug-cc-pVTZ basis set.
@@ V Conclusions
A large number of recent computational studies have explored the influence of strong electron–photon coupling on the electronic structure of cavity-embedded molecular systems. Many such studies have focused on the impact that strong coupling can have on ground-state properties, with comparatively fewer works focusing on molecular response properties, particularly higher-than-first-order responses. In this work, we have developed mean-field approaches to model the response of a molecule within an optical cavity to static external electric fields. While the electric dipole polarizability (a first-order response property) is not particularly sensitive to the presence of the cavity, the first hyperpolarizability (a second-order property) can be significantly impacted by the cavity, depending on the cavity mode polarization axis. Our calculations also indicate that not only are first hyperpolarizabilities more sensitive to the basis set size and composition when compared to polarizabilities, which is well known (see, e.g., Ref. 135, 136, 137, 138), but the cavity-induced changes in are also more affected by the basis set choice than the corresponding changes in . Interestingly, even though the static polarizabilities are less sensitive to basis sets than the hyperpolarizabilities, the use of small basis set and the lack of diffuse functions can still lead to qualitatively incorrect predictions of cavity-induced changes to as shown in Fig. 3 (A). Moving forward, it will be interesting to examine cavity-induced changes to higher-order static response properties, as well as their dynamic analogues, in which case we expect the cavity mode frequency to play an important part. Lastly, it also will be important to consider how explicit electron–photon correlations that are captured by many-body ab initio QED approaches impact the response properties of cavity-embedded systems, as it has repeatedly been demonstrated that mean-field approaches such as those studied herein tend to exaggerate cavity-induced changes to electronic structure.[69, 54, 55]
@@
The cartesian coordinates for -nitroaniline obtained from geometry optimizations using HF, SVWN, PBE, B3LYP, and B97X and the d-aug-cc-pVTZ basis set, the ground-state energies and response properties of -nitroaniline computed using QED-HF and QED-DFT at different values and a fixed cavity frequency of 0.1 E, the excitation energies and oscillator strenghts of the first bright state of -nitroaniline obtained using QED-HF and QED-DFT, the ground-state energies and response properties of -nitroaniline computed using QED-HF and QED-DFT at a.u. and different cavity frequencies, and the response properties of -nitroaniline computed using HF at different values and basis sets with a fixed cavity frequency of 0.1 E.
Acknowledgements